# GraphLoRA: Empowering LLMs Fine-Tuning via Graph Collaboration of MoE

### v1

@@ 5.4 Conclusion
In this paper, we address the instability problem of LLMs due to the imbalance load issue of MoE. We propose a novel framework GraphLoRA to enhance the collaborations of multiple experts in parameter efficient fine-tuning of LLMs. GraphLoRA empowers the graph router to assign weights to experts by effective information sharing among experts on graph neural networks. Besides, two novel coordination strategies, i.e., the Poisson distribution-based expert distinction strategy and the normal distribution-based load balance strategy are proposed in our work. The two constraint operations further enhance the capability of each expert and comprehensively coordinate the collaboration among them. The efficient collaboration among our sparse MoE architecture provides a solution to make trade-offs among the model performance, stability, and resource overhead in the literature of LLMs. Due to the limitation of computation resources, we conducted experiments to verify the effectiveness of our method with different LLMs under 10B parameters. In future work, we will attempt to conduct experiments on more powerful LLMs with larger parameter sizes.

### v3

@@ 5.5 Conclusion
We propose GMoE, a novel framework to address the instability of LLMs caused by MoE load imbalance. GMoE enhances expert collaboration in parameter-efficient fine-tuning through effective information sharing on graph neural networks. GMoE introduces a novel Poisson-based expert distinction strategy to promote expert specialization while employing a normal-based load balance strategy to regulate workload distribution. The comprehensive collaboration among our sparse graph MoE architecture provides a solution to make trade-offs among the model performance, stability, and resource overhead.
5.6 Limitation
We propose GMoE, a graph-based Mixture-of-Experts framework that addresses load imbalance and instability in LLM fine-tuning by enabling expert collaboration via a GNN-powered graph router. Despite its advancements, GMoE has several notable limitations: (1) due to computational resource constraints, the framework’s effectiveness and multi-expert balance mechanism have only been validated in downstream task fine-tuning for LLMs with under 10B parameters, leaving their applicability to larger-scale pre-training settings (e.g., 100B+ parameters) untested; (2) while initial random expert connections in the MoE Graph show minimal impact on performance, the current static graph topology does not adaptively optimize expert interactions, potentially limiting the full realization of collaborative potential and requiring future exploration of dynamic structure learning. These gaps highlight opportunities for future research to scale GMoE to larger models, dynamic graph architectures, and expanded application domains.